I AM LOST THIS CHAPTER PLEASE HELP ASAP

Lane Company manufactures a single product and applies overhead cost to that product using standard direct labor-hours. The bRequired: 1. Compute the predetermined overhead rate for the year. Break the rate down into variable and fixed elements. 2. Peyuncu. - Compute the predetermined overhead rate for the year. Break the rate down into variable and fixed elements. 2. Prep1. Compute the predetermined overhead rate for the year. Break the rate down into variable and fixed elements. 2. Prepare a s1. Compute the predetermined overhead rate for the year. Break the rate down into variable and fixed elements 2. Prepare a st
Answer #1

Solution 1 and 2:

G Н 2 3 4 5 Solution 1: 6 Computation of Predetermined overhead rate - Lane Company 7 Particulars Amount Budgeted variable ma

Solution 3a:

Standard hours allowed = 156000*1.5 = 234,000 hours

Solution 3b:

R 4 5 6 7 solution 3b: Manufacturing Overhead 10 Particulars Debit Particulars Credit $2,901,600 Applied overhead(234000*$12.

Solution 4:

Variable overhead actual rate = $633750/ 253500 = $2.50 per hour

Variable overhead rate variance = (SP - AP) *Actual Hours = ($4.20 - $2.50) * 253500 = $430,950 Favorable

Variable overhead Efficiency variance = (Standard hours - Actual Hours) *SP = (234000 - 253500)*$4.20 = $81,900 (Unfavorable)

Fixed overhead budget Variance = Budgeted Fixed Overhead - Actual Fixed overhead = $1599000 - $1774500 = $175,500 (unfavorable)

Fixed Overhead Volume Variance = ($8.20*234000) - $1599000 = $319,800 Favorable